# Feedback control of twisted states in the Kuramoto model on nearest neighbor and complete simple graphs

### v1

@@ 4.3. Case of $\sigma\neq 0$
Here by the translation symmetry, Eq.  (4.8) must depend only on $r$ and $\mu$, even if the higher-order terms are included. We take such a sufficiently large value for $b_{3}$ as $\beta_{\sigma}>0$. Noting that $\nu_{q}>0$ for $\kappa>0$ by Proposition 3.3, we easily show that a Hopf bifurcation [26, 29, 35] occurs in (4.8) as follows (cf. Fig. 2):
@@
    As in Remark 4.3(iii), the twisted solution (1.11) ((resp. $\tilde{\mathscr{U}}^{q})$ behaves as if it is an asymptotically stable or unstable solution ((resp. a asymptotically stable families of solutions)) in the KM (1.1) near $b_{1}=b_{1q}$ for $n>0$ sufficiently large. Thus, the KM (1.1) suffers a “bifurcation” similar to one detected in Theorem 4.4 for the CL (1.10).
@@ 5. Numerical Simulations: Nearest Neighbor Graphs
We carried out numerical simulations for the KM (1.1), using the DOP853 solver [28], for $q\in[4]$. We took $n=1000$ and chose the initial values $u_{k}^{n}(0)$, $k\in[n]$, independently randomly according to the uniform distribution on the intervals $\pi+2\pi qk/n,\pi+2\pi qk/n]$ centered at the $q$-twisted state (1.6) with $\Omega_{\mathrm{D}}=0$. So if there is an asymptotically stable that are different from the twisted and modulated or oscillating twisted states, then the responses of (1.1) may converge to it as $t\to\infty$. We also considered two cases $\sigma=0$ and $\pi/3$ for the phase lag, and $\kappa=0.4$ and $0.5$ for the neighbor size. Recall that the $\kappa$-nearest neighbor graph reduce to a complete simple one when $\kappa=0.5$.
@@
In Figs. 7 and 8, $u_{k}^{n}(t)$, $k\in[n]$, at $t=1000$, which may be regarded as the steady states from Figs. 5 and 6, are plotted as small red disks for $\sigma=0$ and $\pi/3$, respectively. Here the same values of $b_{1}$ and $u_{k}^{n}(0)$, $k\in[n]$, as in Figs. 5 and 6 were used. We observe that the responses of the KM (1.1) converge to the twisted and modulated or oscillating twisted states, respectively, for the larger and smaller values of $b_{1}$, as predicted by Theorems 4.2 and 4.4 with the assistance of Corollary 2.5 and Theorem 2.8. Indeed, we confirmed that the deviation from the twisted state is about $10^{-7}$ at most in the left column of each figure for the larger values of $b_{1}$. In particular, the target state (1.6) is accomplished there.
@@
In Figs. 9 and 10, the deviation, $u_{k}^{n}(t)-2\pi qk/n$, $k\in[n]$, of the steady state in the right columns of Figs. 7 and 8 from the $q$-twisted state (1.6) in the KM (1.1) for $\sigma=0$ and $\sigma=\pi/3$, respectively, when $b_{1}$ is considered to be smaller than the bifurcation point, is plotted as small red disks. It was also estimated from the most probably leading term displayed in Figs. 7 and 8 and is plotted as a blue line. The agreement between both results is fine except in Fig. 9(a) for $q=1$ and $\kappa=0.4$. The reason for their disagreement in Fig. 9(a) is considered to be that the absolute value of $\mu_{2q}=(\chi_{1}(2q,q)-\chi_{1}(q,q))\cos\sigma$ is small and the $2q$-oscillation mode is easily to be excited when $q=1$ and $\kappa=0.4$ (see Fig. 1), compared with the other cases.
@@
Finally, we present numerically computed bifurcation diagrams for $\sigma=0$ and $\pi/3$ in Figs. 11 and 12, respectively. The amplitude $r$ of the expression (5.2) estimated from the numerical simulation results for the steady states as in Figs. 7-10 are plotted as small red disks, and the theoretical predictions,
@@ 6. Numerical Simulations: Complete Simple Graphs
Figures 13 and 14 show the time-histories of every $100$th node (from 50th to 950th) for $\sigma=0$ and $\pi/3$, respectively, like Figs. 5 and 6. Here the values of $u_{k}^{n}(t)\mod 2\pi$, $k\in[n]$, are plotted as the ordinates. The values of $b_{1}=0.52$ and $0.48$ (resp. $b_{1}=0.26$ and $0.24$) were chosen in the left and right columns of Fig. 13 (resp. of Fig. 14), respectively, and they are lager and smaller than the bifurcation point approximated by $b_{1q}=0.5$ (resp. $b_{1q}=0.25$) (see Table. 1). We see that the responses converge to the steady states rapidly, and oscillations occur in the right column of Fig. 14 for $\sigma=\pi/3$ and $b_{1}=0.24$, as detected by Theorem 4.4 for the CL (1.10).
In Figs. 15 and 16, $u_{k}^{n}(t)$, $k\in[n]$, at $t=1000$, which may be regarded as the steady states from the results of Figs. 13 and 14, are plotted as small red disks for $\sigma=0$ and $\pi/3$, respectively. Here the same values of $b_{1}$ and $u_{k}^{n}(0)$, $k\in[n]$, as in Figs. 13 and 14 were used. We observe that the responses of the KM (1.1) converge to the twisted and modulated or oscillating twisted states for $b_{1}=0.52$ or $0.26$ and $0.48$ or $0.24$ in the left and right columns of each figure, respectively, as predicted by Theorems 4.2 and 4.4 with the assistance of Corollary 2.5 and Theorem 2.8. Indeed, we confirmed that the deviation from the twisted state is about $10^{-11}$ and $10^{-6}$ at most in the left columns of Figs. 15 and 16 for $b_{1}=0.52$ and $0.26$, respectively. In particular, the target state (1.6) is accomplished there. The most probably leading term (5.2) in the modulated and oscillating twisted solutions (4.7) and (4.12) was also estimated from the numerical simulation results by using the least mean square method and is plotted as a blue line in each figure, as in Figs. 7 and 8. Both results coincide almost completely, as detected by Theorems 4.2 and 4.4 for the CL (1.10).
In Figs. 17 and 18, the deviation, $u_{k}^{n}(t)-2\pi qk/n$, $k\in[n]$, of the steady state in the right columns of Figs. 15 and 16 from the desired $q$-twisted one in the KM (1.1) with $\sigma=0$ and $\sigma=\pi/3$, respectively, is plotted as small red disks. Estimates obtained from the most probably leading terms displayed in Figs. 15 and 16 are also plotted as blue lines, as in Figs. 9 and 10. Both results coincide almost completely.
Finally, we present numerically computed bifurcation diagrams for $\sigma=0$ and $\pi/3$ in Figs. 19 and 20, respectively, as in Figs. 11 and 12. The amplitude $r$ of (5.2) estimated from the numerical simulation results are plotted as small red disks, and the theoretical predictions given by (5.4) are plotted as black solid lines. Good agreement between both results is found, especially in Fig. 19, although slight differences are seen in Fig. 20.
@@ 7. Concluding Remarks
We summarize this paper as follows: We studied feedback control of twisted states in the KM (1.1) of identical oscillators defined on deterministic nearest neighbor graphs containing complete simple ones when it may have phase-lag. Using the center manifold reduction techniqye [29], we analyzed the stability and bifurcations of twisted solutions in the CL (1.10) for the KM (1.1) subjected to feedback control. In particular, it was shown that the twisted solutions exist and can be stabilized not only for nearest neighbor graphs but also for complete simple graphs. Moreover, the CL (1.10) was shown to suffer bifurcations at which the twisted solutions becomes unstable and a stable one-parameter family of modulated or oscillating twisted solutions is born, depending on whether the phase-lag is zero or not. We demonstrated the theoretical results by numerical simulations for the feedback controlled KM (1.1) on deterministic nearest neighbor and complete simple graphs.
Finally, we give some comments for future work. The KM and its generalization with phase-lag was studied for different solutions from twisted ones in [51, 11, 50, 5, 44, 47, 37]. In particular, chimera states were discussed in [50, 5, 44] The theory reviewed in Section 2 is also applicable to these cases and may be useful to uncover their dynamics. The KM with time delay, which possesses important applications in neuroscience [14, 9, 20], has often been a subject of research [1, 4, 52, 55, 47]. It will be another next target to extend the theory of Section 2 to the KM with time delay and its CL.
@@ Acknowledgements
$\qquad\quad+(\xi_{q}\eta_{j}+\xi_{j}\eta_{1})\sin 2\pi(q+j)x)\sin 2\pi qx\bigr)+\cdots.\hskip 80.00012pt$

### v2

@@ 4.3. Case of $\sigma\neq 0$
Here by the translation symmetry, Eq.  (4.8) must depend only on $r$ and $\mu$, even if the higher-order terms are included, like (4.5). We take such a sufficiently large value for $b_{3}$ as $\beta_{\sigma}>0$. Noting that $\nu_{q}>0$ for $\kappa>0$ by Proposition 3.3, we easily show that a Hopf bifurcation [26, 29, 35] occurs in (4.8) as follows (cf. Fig. 2):
@@
    As in Remark 4.3(iii), the target orbit (1.3) ((resp. $\tilde{\mathscr{U}}^{q})$ behaves as if it is an asymptotically stable solution or it is actually unstable ((resp. as if it is an asymptotically stable families of solutions)) in the KM (1.1) near $b_{1}=b_{1q}$ for $n>0$ sufficiently large. Thus, the KM (1.1) suffers a “bifurcation” similar to one detected in Theorem 4.4 for the CL (1.6).
@@ 5. Numerical Simulations: Nearest Neighbor Graphs
We carried out numerical simulations for the KM (1.1), using the DOP853 solver [28], for $q\in[4]$. We took $n=1000$ and chose the initial values $u_{k}^{n}(0)$, $k\in[n]$, independently randomly according to the uniform distribution on the intervals $\pi+2\pi qk/n,\pi+2\pi qk/n]$ centered at the $q$-twisted state (1.3) with $\Omega_{\mathrm{D}}=0$. So if there is an asymptotically stable that are different from the twisted and modulated or oscillating twisted states, then the responses of (1.1) may converge to it as $t\to\infty$. We also considered two cases $\sigma=0$ and $\pi/3$ for the phase lag, and $\kappa=0.4$ and $0.5$ for the neighbor size. Recall that the $\kappa$-nearest neighbor graph reduces to a complete simple one when $\kappa=0.5$.
@@
Figure 7. Space-time plots of all oscillator phases $u_{k}^{n}(t)$, $k\in[n]$, in the KM (1.1) with $n=1000$, $\kappa=0.4$ and $\sigma=0$: (a) $(q,b_{1},b_{3})=(1,0.16,1)$; (b) $(1,0.12,1)$; (c) $(2,0.275,0.5)$; (d) $(2,0.255,0.5)$; (e) $(3,0.17,0.5)$; (f) $(3,0.15,0.5)$; (g) $(4,0.245,0.5)$; (h) $(4,0.225,0.5)$.
Figure 8. Space-time plots of all oscillator phases $u_{k}^{n}(t)$, $k\in[n]$, in the KM (1.1) with $n=1000$, $\kappa=0.4$, $\sigma=\pi/3$ and $b_{3}=0.5$: (a) $(q,b_{1})=(1,0.08)$; (b) $(1,0.06)$; (c) $(2,0.275)$; (d) $(2,0.255)$; (e) $(3,0.17)$; (f) $(3,0.15)$; (g) $(4,0.245)$; (h) $(4,0.225)$. See also the caption of Fig. 5.
Figures 7 and 8 show space-time plots of all oscillator phases $u_{k}^{n}(t)$, $k\in[n]$, for $\sigma=0$ and $\pi/3$, respectively. Here the same values of $b_{1}$ and $u_{k}^{n}(0)$, $k\in[n]$, as in Figs. 5 and 6 were used. We see that all phases of the KM (1.1) converge to their steady states rapidly, as in Figs. 5 and 6, and that they exhibit oscillations for the smaller values of $b_{1}$ when $\sigma=\pi/3$, in the right column of Fig. 8, whereas only small differences between the larger and smaller values of $b_{1}$ are observed in Fig. 7. Moreover, in both Figs. 7 and 8, the variation in $u_{k}^{n}(t)$ from $k=1$ to $n$ increases as $q$ increases.
@@
In Figs. 9 and 10, $u_{k}^{n}(t)$, $k\in[n]$, at $t=1000$, which may be regarded as the steady states from Figs. 5 and 6, are plotted as small red disks for $\sigma=0$ and $\pi/3$, respectively. Here the same values of $b_{1}$ and $u_{k}^{n}(0)$, $k\in[n]$, as in Figs. 5 and 6 were used. We observe that the responses of the KM (1.1) converge to the twisted and modulated or oscillating twisted states, respectively, for the larger and smaller values of $b_{1}$, as predicted by Theorems 4.2 and 4.4 with the assistance of Corollary 2.5 and Theorem 2.8. Indeed, we confirmed that the deviation from the twisted state is about $10^{-7}$ at most in the left column of each figure for the larger values of $b_{1}$. In particular, the target state (1.3) is accomplished there.
@@
In Figs. 11 and 12, the deviation, $u_{k}^{n}(t)-2\pi qk/n$, $k\in[n]$, of the steady state in the right columns of Figs. 9 and 10 from the $q$-twisted state (1.3) in the KM (1.1) for $\sigma=0$ and $\sigma=\pi/3$, respectively, when $b_{1}$ is considered to be smaller than the bifurcation point, is plotted as small red disks. It was also estimated from the most probably leading term displayed in Figs. 9 and 10 and is plotted as a blue line. The agreement between both results is fine except in Fig. 11(a) for $q=1$ and $\kappa=0.4$. The reason for their disagreement in Fig. 11(a) is considered to be that the absolute value of $\mu_{2q}=(\chi_{1}(2q,q)-\chi_{1}(q,q))\cos\sigma$ is small and the $2q$-oscillation mode is easily to be excited when $q=1$ and $\kappa=0.4$ (see Fig. 1), compared with the other cases.
@@
Finally, we present numerically computed bifurcation diagrams for $\sigma=0$ and $\pi/3$ in Figs. 13 and 14, respectively. The amplitude $r$ of the expression (5.2) estimated from the numerical simulation results for the steady states as in Figs. 9-12 are plotted as small red disks, and the theoretical predictions,
@@ 6. Numerical Simulations: Complete Simple Graphs
Figures 15 and 16 show the time-histories of every $100$th node (from 50th to 950th) for $\sigma=0$ and $\pi/3$, respectively, like Figs. 5 and 6. Here the values of $u_{k}^{n}(t)\mod 2\pi$, $k\in[n]$, are plotted as the ordinates. The values of $b_{1}=0.52$ and $0.48$ (resp. $b_{1}=0.26$ and $0.24$) were chosen in the left and right columns of Fig. 15 (resp. Fig. 16), respectively, and they are larger and smaller than the bifurcation point approximated by $b_{1q}=0.5$ (resp. $b_{1q}=0.25$) (see Table. 1). We see that the responses converge to the steady states rapidly, and oscillations occur in the right column of Fig. 16 for $\sigma=\pi/3$ and $b_{1}=0.24$, as detected by Theorem 4.4 for the CL (1.6).
In Figs. 17 and 18, $u_{k}^{n}(t)$, $k\in[n]$, at $t=1000$, which may be regarded as the steady states from the results of Figs. 15 and 16, are plotted as small red disks for $\sigma=0$ and $\pi/3$, respectively. Here the same values of $b_{1}$ and $u_{k}^{n}(0)$, $k\in[n]$, as in Figs. 15 and 16 were used. We observe that the responses of the KM (1.1) converge to the twisted and modulated or oscillating twisted states for $b_{1}=0.52$ or $0.26$ and $0.48$ or $0.24$ in the left and right columns of each figure, respectively, as predicted by Theorems 4.2 and 4.4 with the assistance of Corollary 2.5 and Theorem 2.8. Indeed, we confirmed that the deviation from the twisted state is about $10^{-11}$ and $10^{-6}$ at most in the left columns of Figs. 17 and 18 for $b_{1}=0.52$ and $0.26$, respectively. In particular, the target state (1.3) is accomplished there. The most probably leading term (5.2) in the modulated and oscillating twisted solutions (4.7) and (4.12) was also estimated from the numerical simulation results by using the least mean square method and is plotted as a blue line in each figure, as in Figs. 9 and 10. Both results coincide almost completely, as detected by Theorems 4.2 and 4.4 for the CL (1.6).
In Figs. 19 and 20, the deviation, $u_{k}^{n}(t)-2\pi qk/n$, $k\in[n]$, of the steady state in the right columns of Figs. 17 and 18 from the desired $q$-twisted one in the KM (1.1) with $\sigma=0$ and $\sigma=\pi/3$, respectively, is plotted as small red disks. Estimates obtained from the most probably leading terms displayed in Figs. 17 and 18 are also plotted as blue lines, as in Figs. 11 and 12. Both results coincide almost completely.
Finally, we present numerically computed bifurcation diagrams for $\sigma=0$ and $\pi/3$ in Figs. 21 and 22, respectively, as in Figs. 13 and 14. The amplitude $r$ of (5.2) estimated from the numerical simulation results are plotted as small red disks, and the theoretical predictions given by (5.4) are plotted as black solid lines. Good agreement between both results is found, especially in Fig. 21, although slight differences are seen in Fig. 22.
@@ 7. Concluding Remarks
We summarize this paper as follows: We studied feedback control of twisted states in the KM (1.1) of identical oscillators defined on deterministic nearest neighbor graphs containing complete simple ones when it may have phase-lag. Using the center manifold reduction technique [29], we analyzed the stability and bifurcations of twisted solutions in the CL (1.6) for the KM (1.1) subjected to feedback control. In particular, it was shown that the twisted solutions exist and can be stabilized not only for nearest neighbor graphs but also for complete simple graphs. Moreover, the CL (1.6) was shown to suffer bifurcations at which the twisted solutions becomes unstable and a stable one-parameter family of modulated or oscillating twisted solutions is born, depending on whether the phase-lag is zero or not. We demonstrated the theoretical results by numerical simulations for the feedback controlled KM (1.1) on deterministic nearest neighbor and complete simple graphs.
From an applied viewpoint, the present work suggests that such a simple feedback law as the linear-cubic terms in (1.1) can be used to realize and maintain prescribed twisted phase patterns in coupled oscillator networks by tuning only a few scalar gains. Kuramoto-type phase descriptions are widely used as reduced models across disciplines, including power-system and microgrid settings where the stability of phase-locked operating states and their control under disturbances are central [22, 15, 27, 2, 58], and biological and neural oscillator networks where spatiotemporal rhythms and network level oscillatory patterns are of interest [8, 5, 61, 72, 73]. For the KM (1.1), the CL analysis provides explicit stability regions and bifurcations of twisted solutions, thereby offering analytically tractable guidance for gain tuning and mode selection. More broadly, the results fit into the general theme of controlling collective behavior in complex networks under practical constraints [13, 48, 17, 59]. Although our main analysis is carried out in the CL framework, it is reasonable to expect that it remains at least qualitatively valid for moderate-size graphs (e.g., $n\sim 50$). The CL predictions can then be used to indicate parameter regions where qualitative transitions occur, which can be tested numerically on such moderate-size graphs.
Finally, we give some comments for future work. The KM and its generalization with phase-lag was studied for different solutions from twisted ones in [51, 11, 50, 5, 44, 47, 37]. In particular, chimera states were discussed in [50, 5, 44]. The theory reviewed in Section 2 is also applicable to these cases and may be useful to uncover their dynamics. The KM with time delay, which possesses important applications in neuroscience [14, 9, 20], has often been a subject of research [1, 4, 52, 55, 47]. It will be another next target to extend the theory of Section 2 to the KM with time delay and its CL.
@@ Acknowledgements
$\qquad\quad+(\xi_{q}\eta_{j}+\xi_{j}\eta_{1})\sin 2\pi(q+j)x)\sin 2\pi qx\bigr)+\cdots.\hskip 100.00015pt$